# Adaptive Extraction Network for Multivariate Long Sequence Time-Series Forecasting

### v1

@@ 4.5 Visualization of forecasting
To evaluate the predictive performance of ATVCNet compared to SOTA MLSTF models (such as ConvTimeNet, TimeMixer, iTransformer, TSMixer and PatchTST) in real-world scenarios, we qualitatively compared the prediction results of the last dimension on the test set of the ETTh1 dataset, as shown in Figure 4. The input length was set to 96, and the prediction length was set to 192 to assess the fitting between the predicted sequences and the actual sequences.
@@
From Figure 4, it can be observed that ATVCNet is capable of adapting to the dynamic changes in time series data, capturing both the periodic and trend information. However, while other models effectively capture the periodic information of the time series, they neglect to acquire trend information. For instance, around the 280th time point, the sequence suddenly oscillates downward. ATVCNet accurately captured this oscillatory trend, while other models retained the trend information from the previous time period, indicating that their predictive performance did not benefit from the module that extracts long-term feature information. In summary, ATVCNet not only captures local context information but also effectively captures the global trend information of the time series, resulting in better fitting results.
@@ 4.6 Long-term information utilization
To validate ATVCNet’s ability to capture long-term correlations, we compared the performance of different models on the ETTh1 and ETTm1 datasets, with output lengths set to $\left\{96,720\right\}$ and input sequence windows set to $\left\{24,48,96,192,336\right\}$. In theory, using longer input sequence windows can increase the receptive field and potentially improve prediction performance. However, as evidenced by the experimental results in Figure 5, the performance of Transformer-based models, TimeMixer, and MICN models does not continuously improve with increasing model input length, indicating that these models do not benefit from longer input sequence windows, i.e., they lack in capturing long-term correlations. In contrast, ATVCNet, ConvTimeNet, and TSMixer show continuous improvement in model performance (decreasing MSE values) with increasing input sequence length, demonstrating that these models can exploit more features from longer input sequence windows and effectively capture the long-term correlations in the input sequence.
@@ 4.7 T-tests
To further validate the effectiveness of the ATVCNet model compared to other baseline models, we conducted T-tests to assess the significant differences between the prediction results of ATVCNet and the baseline models. As shown in Figure 6, we can draw the following conclusions: (1) The T-statistic for ATVCNet relative to all baseline models significantly exceeds $T^{\prime}=1.782$ (by looking up the critical value in the T-distribution table based on the P-value, we obtain the T-value), indicating that the average MSE of ATVCNet is significantly lower than that of the baseline models. (2) The calculated P-values are all less than the threshold value of $P=0.05$, indicating statistically significant differences in prediction results among different models. Based on the above analysis, we can conclude that the observed performance differences between ATVCNet and the baseline models are indeed the result of true differences, rather than random factors.
Figure 6: The result of the statistical significance test between ATVCNet and fourteen baseline models.
@@ 4.8 Efficiency analysis
To evaluate the efficiency of ATVCNet, we selected the ETTh1 dataset and compared it with six SOTA models. We conducted efficiency analysis of the models from four aspects: computational complexity (Flops), model size (Parameters), training time, and inference time. To facilitate a clearer comparison and analysis of model efficiency, we integrated these four metrics and computed their average values as the model efficiency scores, as shown in Table 3. Table 3 distinctly demonstrates the efficiency advantages of ATVCNet, which are particularly crucial for practical applications of MLSTF. This advantage mainly stems from the model’s parameter sharing mechanism, deformable convolutions, and the ability to rapidly update network weights through pseudo-inverse algorithms. Additionally, we observed a decreasing trend in the training time of iTransformer with increasing input length. This is because iTransformer employs an early termination strategy during training, where training halts when the loss value decreases continuously for a certain number of iterations (n=3 in iTransformer). According to our experiments, as the input length increases, the model is more likely to meet this condition during training.
@@ 4.9 Ablation study
To assess the effectiveness of each component within ATVCNet, we conducted the following ablations: (1) w/o Variable: removal of modules designed for extracting inter-variable correlation information; (2) w/o Temporal: elimination of modules for extracting temporal correlation information; (3) w/o ALL: complete removal of all feature extraction modules. We conducted ablation studies on the ETTh2 and Electricity datasets, with input lengths set to 96 and prediction lengths set to $\left\{96,192,336,720\right\}$.
The results in Table 4 indicate that removing any module from the original model led to a decrease in performance. Specifically: (1) Removing the variable module (w/o Variable) results in the model losing its ability to extract interdependencies among different variables. By solely relying on the the time-domain information extracted by the model, the prediction accuracy of the model is reduced. (2) Removing the temporal domain feature extraction module (w/o Temporal) results in the model’s inability to extract long-term trend information and long-term correlations from the time series. Relying solely on the information extracted among different variables by the model also leads to a decline in the model’s prediction performance. (3) Completely removing the feature extraction module (w/o ALL) is equivalent to using only the FFN layer for prediction. Experimental results indicate that this approach performs poorly, once again demonstrating that FFN lacks the ability to effectively extract complex features from time series data.
@@
Models ATVCNet w/o Variable w/o Temporal w/o All
@@ 5 Conclusion
This paper introduces ATVCNet, an adaptive temporal-variable convolutional network. From both the temporal and variable dimensions, we utilize convolutional kernels with different dilation factors to capture underlying local patterns and features at different resolutions within complex time series. Additionally, we employ adaptive average pooling to learn global temporal patterns. Leveraging the flexibility of deformable convolutions, the network can effectively capture correlations and dynamics between different variables. Furthermore, considering the inherent variability in the data distribution of time series, the model allows for dynamic updates based on changes in the data. Through extensive experimentation on eight benchmark datasets, our findings demonstrate the superior performance of ATVCNet compared to previous SOTA models.

### v2

@@ 4.5 Visualization of forecasting
To evaluate the predictive performance of ACNet compared to SOTA TSF models (such as ConvTimeNet, TimeMixer, iTransformer, TSMixer and PatchTST) in real-world scenarios, we qualitatively compared the prediction results of the last dimension on the test set of the ETTh1 dataset, as shown in Figure 5. The input length was set to 96, and the prediction length was set to 192 to assess the fitting between the predicted sequences and the actual sequences.
@@
From Figure 5, it can be observed that ACNet is capable of adapting to the dynamic changes in time series data, capturing both the periodic and trend information. However, while other models effectively capture the periodic information of the time series, they neglect to acquire trend information. For instance, around the 280th time point, the sequence suddenly oscillates downward. ACNet successfully captured the trend information of such non-stationary oscillatory changes, while other models retained the trend information from the previous time period, indicating that their predictive performance did not benefit from the module that extracts long-term feature information. In summary, ACNet can not only capture local contextual information but also effectively capture the global trend information and nonlinear information of time series, thereby achieving better fitting results.
@@ 4.6 Long-term information utilization
To validate ACNet’s ability to capture long-term correlations, we compared the performance of different models on the ETTh1 and ETTm1 datasets, with output lengths set to $\left\{96,720\right\}$ and input sequence windows set to $\left\{24,48,96,192,336\right\}$. In theory, using longer input sequence windows can increase the receptive field and potentially improve prediction performance. However, as evidenced by the experimental results in Figure 6, the performance of Transformer-based models, TimeMixer, and MICN models does not continuously improve with increasing model input length, indicating that these models do not benefit from longer input sequence windows, i.e., they lack in capturing long-term correlations. In contrast, ACNet, ConvTimeNet, and TSMixer show continuous improvement in model performance (decreasing MSE values) with increasing input sequence length, demonstrating that these models can exploit more features from longer input sequence windows and effectively capture the long-term correlations in the input sequence.
@@ 4.7 T-tests
To further validate the effectiveness of the ACNet model compared to other baseline models, we conducted T-tests to assess the significant differences between the prediction results of ACNet and the baseline models. As shown in Figure 7, we can draw the following conclusions: (1) The T-statistic for ACNet relative to all baseline models significantly exceeds $T^{\prime}=1.782$ (by looking up the critical value in the T-distribution table based on the P-value, we obtain the T-value), indicating that the average MSE of ACNet is significantly lower than that of the baseline models. (2) The calculated P-values are all less than the threshold value of $P=0.05$, indicating statistically significant differences in prediction results among different models. Based on the above analysis, we can conclude that the observed performance differences between ACNet and the baseline models are indeed the result of true differences, rather than random factors.
Figure 7: The result of the statistical significance test between ACNet and fourteen baseline models.
4.8 Validation of capturing nonlinear features
To further verify the effectiveness of our proposed nonlinear feature adaptive extraction module in extracting features from real-world time series data, we selected the Exchange dataset for validation. We processed the data using both standard convolutional kernels and our proposed gated deformable convolution (GDC). We used a batch of the dataset and extracted feature maps from the intermediate layers to observe the activation of input data after passing through the convolutional layers. This helps in understanding how different convolutional modules capture nonlinear relationships. We visualized the feature maps from the first eight layers, as shown in Figure 8:
(a) Standard Converlutional Feature Map.
(b) GDC Feature Map.
Figure 8: Visualization of feature maps from different convolutions.
From Figure 8, we can draw the following conclusions:
(1) Amplitude of Variations in Feature Maps. The feature maps from standard convolution exhibit smaller amplitudes of change with fewer fluctuations, indicating a potential inadequacy in handling sudden and nonlinear changes in the input data. The feature maps from GDC exhibit larger amplitudes of change with more noticeable fluctuations, indicating a greater flexibility in capturing sudden and nonlinear changes in the input data.
(2) Diversity in Feature Maps. The feature maps from standard convolution show more consistent patterns with relatively smooth changes. The feature maps from GDC show more diverse patterns and variations, indicating a stronger adaptability in handling nonlinear features.
(3) Response to Sudden Changes. The feature maps from standard convolution exhibit smooth transitions at points of sudden changes in the data, lacking sensitivity to such changes. The feature maps from GDC exhibit noticeable fluctuations and variations at points of sudden changes in the data, better capturing these sudden features.
Specifically, from the standard convolutional feature maps (as shown in Figure 8(a)), we can see that most feature maps show relatively smooth and stable changes with almost no significant fluctuations or sudden changes. In feature map 5, some fluctuations can be observed, but the overall changes remain relatively stable. From the GDC feature maps (as shown in Figure 8(b)), we can see that feature maps 1, 3, 4, and 5 exhibit significant fluctuations and sudden changes, indicating that GDC can capture dynamic sudden features in the input data. Feature maps 0 and 2 exhibit more details and fluctuations, suggesting that GDC performs well in handling non-stationary features.
In summary, GDC performs better in handling nonlinear features because it can dynamically adjust the positions of the convolutional kernels, making it more adaptable to variations in the input data.
@@ 4.9 Efficiency analysis
To evaluate the efficiency of ACNet, we selected the ETTh1 dataset and compared it with six SOTA models. We conducted efficiency analysis of the models from four aspects: computational complexity (Flops), model size (Parameters), training time, and inference time. To facilitate a clearer comparison and analysis of model efficiency, we integrated these four metrics and computed their average values as the model efficiency scores, as shown in Table 4. Table 4 distinctly demonstrates the efficiency advantages of ACNet, which are particularly crucial for practical applications of TSF. This advantage mainly stems from the model’s parameter sharing mechanism, deformable convolutions, and the ability to rapidly update network weights through pseudo-inverse algorithms. Additionally, we observed a decreasing trend in the training time of iTransformer with increasing input length. This is because iTransformer employs an early termination strategy during training, where training halts when the loss value decreases continuously for a certain number of iterations (n=3 in iTransformer). According to our experiments, as the input length increases, the model is more likely to meet this condition during training.
@@ 4.10 Ablation study
To assess the effectiveness of each component within ACNet, we conducted the following ablations: (1) w/o GDC: remove the gated deformable convolution module used for extracting nonlinear features of time series; (2) w/o Temporal: elimination of modules for extracting temporal correlation information; (3) w/o ALL: complete removal of all feature extraction modules. We conducted ablation studies on the ETTh2 and Electricity datasets, with input lengths set to 96 and prediction lengths set to $\left\{96,192,336,720\right\}$.
The results in Table 5 indicate that removing any module from the original model led to a decrease in performance. Specifically: (1) Removing the GDC module (w/o GDC) will cause the model to lose its ability to extract nonlinear information from the time series. By solely relying on the the time-domain information extracted by the model, the prediction accuracy of the model is reduced. (2) Removing the temporal feature extraction module (w/o temporal) will prevent the model from extracting local contextual information and global patterns from the time series, leading to a decline in predictive performance. (3) Completely removing the feature extraction module (w/o ALL) is equivalent to using only the FFN layer for prediction. Experimental results indicate that this approach performs poorly, once again demonstrating that FFN lacks the ability to effectively extract complex features from time series data.
@@
Models ACNet w/o GDC w/o Temporal w/o All
@@ 5 Conclusion
This paper introduces a feature-driven time series prediction network, ACNet. From the perspective of effective feature extraction and utilization, we use convolutional kernels with different dilation factors to capture latent local patterns and features at different resolutions in complex time series. Additionally, adaptive average pooling is employed to learn the global temporal patterns of the time series. By utilizing improved gated mechanism-based deformable convolutions, the network can effectively capture nonlinear features and pattern information in the time series. Considering the inherent variability in the data distribution of time series, the model allows for dynamic updates based on changes in the data. Through extensive experimentation on twelve benchmark datasets, our findings demonstrate the superior performance of ACNet compared to previous SOTA models.